A school ordered $600 worth of lightbulbs. Some of the lightbulbs cost $1 each and the others cost $2 each. If twice as many $1 bulbs as $2 bulbs were ordered, how many lightbulbs were ordered altogehter?
Answer by checkley75(3666)

2*1*X+1*2*X=600
2X+2X=600
4X=600
X=600/4
X=150 ANSWER FOR THE NUMBER IF $2.00 BULBS.
X=2*150=300 ANSWER FOR THE NUMBER OF $1.00 BULBS.
PROOF
2*150+1*300=600
300+300=600
600=600